# Rapidly Converging Time-Discounted Ergodicity on Graphs for Active Inspection of Confined Spaces

### v1

@@ IV-E Nonuniform Target Distribution
Let $P$ be required to satisfy the detailed balance condition in (38) for a given target distribution $\bar{\rho}$ with no zero entries. Then, the normalized ergodicity deviation in (18) with weighting $Q=\Pi^{-1}$ is minimized by selecting the stochastic matrix $P$ to have the smallest, second-largest eigenvalue,
$\arg\min_{P}\lambda_{2}(P).$ (39)
Additionally, the optimal stochastic matrix $P$ can be obtained by solving the following convex optimization problem
@@
s.t. $\quad\textbf{1}^{T}P=\textbf{1}^{T}$ (Stochastic $(\text{Stochastic }P)$
[…]
$\quad P_{i,j}\geq 0$ (Stochastic $(\text{Stochastic }P)$
$\quad P_{i,j}=0\quad\text{if}\quad(j,i)\notin\mathcal{E}$ (Transitions ${(\mbox{Transitions in}}~{}\mathcal{G})$
@@
$\max_{\rho_{0}}\frac{\|Q^{1/2}(\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho})\|_{2}}{\|Q^{1/2}\left(\rho_{0}-\bar{\rho}\right)\|_{2}}$ (41)
$=\max_{\rho_{0}}\frac{\left\|Q^{1/2}f_{w}\left(P|_{u_{1}^{\perp}}\right)\rho_{0{\perp}}\right\|_{2}}{\|Q^{1/2}\rho_{0\perp}\|_{2}}$
$=\max_{\rho_{0}}\frac{\left\|Q^{1/2}f_{w}\left(P|_{u_{1}^{\perp}}\right)Q^{-1/2}Q^{1/2}\rho_{0{\perp}}\right\|_{2}}{\|Q^{1/2}\rho_{0\perp}\|_{2}}$
$=\max_{\rho_{0}}\frac{\left\|f_{w}\left(Q^{1/2}P|_{u_{1}^{\perp}}Q^{-1/2}\right)Q^{1/2}\rho_{0{\perp}}\right\|_{2}}{\|Q^{1/2}\rho_{0\perp}\|_{2}}$
$=\max_{\rho_{0}}\frac{\left\|f_{w}\left(\tilde{P}|_{u_{1}^{\perp}}\right)\tilde{\rho}_{0{\perp}}\right\|_{2}}{\|\tilde{\rho}_{0\perp}\|_{2}}$
@@
where $\Pi=\text{diag}(\bar{\rho})$ is invertible and its square root exists since $\bar{\rho}$ is strictly positive. Note that $\tilde{P}$ is symmetric from the detailed balance condition in (38), because
$P\Pi$ $=\Pi P^{T}$ (43)
$\Pi^{-1/2}P\Pi^{1/2}$ $=\Pi^{1/2}P^{T}\Pi^{-1/2}$
$\tilde{P}$ $=\tilde{P}^{T},$
[…]
@@ IV-F Convex Bounding for General Markov Chain
The detailed balance condition in the previous subsection can be restrictive for ergodicity. An obvious example is when transitions are allowed in only one direction between some regions, because the detailed balance condition leads to zero probability even for the available transition. Thus, in such one-way transition cases, the detailed balance condition results in larger normalized ergodicity deviation when compared to the case without the condition, as illustrated in Fig. 4.
@@
The detailed balance requirement can be relaxed, for general directed graphs (e.g., with potential one-way transitions), by optimizing an upper bound on the normalized ergodicity deviation, as shown in the theorem below.
@@
Corollary 3.1.
[…]
@@ IV-G Connection to Fastest Mixing Markov Chain
This is equivalent to the fastest mixing Markov chain (FMMC) formulation, where the stochastic matrix $P$ is selected such that the second largest eigenvalue modulus (SLEM), $|\lambda_{2}|$, is minimized. The FMMC formulation upper-bounds the REMC by ensuring fast convergence on the discounted time average since,
@@
The difference in ergodicity achieved with the FMMC formulation that maximizes the second largest eigenvalue modulus (SLEM) and the proposed approach that maximizes the second largest eigenvalue is illustrated in the following example.
@@ V Estimation Framework
We now discuss the estimation module in Fig. 1 that comprises SLAM and anomaly detection. In addition to the occupancy grid map and pose estimates for navigation, the estimator also provides the measure $\mu$ (defined in (3)) to the REMC method. Specifically, this measure is selected as information entropy using a Bayesian anomaly (FOD) detection method. The core idea of detecting FODs is to evaluate the hypothesis that the difference between the local point cloud given by the robot’s depth camera and the reference model of the corresponding section of the confined space is significant. Consequently, the FOD detection method uses the marginal posterior of our SLAM formulation described below.
@@ V-A Hierarchical SLAM
In general, SLAM provides optimal point estimates of robot poses and map features along with representations of the corresponding uncertainties. We specifically use the Graph-SLAM algorithm [47, ch.11.3, p.346] due to its ability to handle non-linear robot motion and sensor models. Graph-SLAM is primarily based on the $\text{SE}(3)$ manifold according to the formulation in [48], and is optimized by the Gauss–Newton algorithm described in [49]. We modify it as a hierarchical SLAM, similar to [50] and [51]. Instead of maintaining the full resolution factor graph and performing hierarchical optimization, marginalization is done as an intermediate EKF-SLAM, and only the sparse factor graph is maintained with the posterior from the EKF-SLAM as the factor between the nodes. In particular, EKF-SLAM is reset and a new pose node is created based on the distance traveled from the most recent pose node. Special nodes, called inspection nodes, are created once the robot reaches the waypoint specified by the waypoint placement method. The corresponding depth point clouds are also stored at these inspection nodes. Furthermore, nodes that fall outside the current estimation horizon are pruned and replaced by a factor node through marginalization. Once the Graph-SLAM is optimized, the estimated poses of all the unpruned nodes, $\mathcal{T}\triangleq\{T_{1},T_{2},\cdots T_{i}\cdots\}$, and the corresponding covariance matrices, $\Sigma_{\mathcal{T}}\triangleq\text{Var}(\mathcal{T})$, are used for recursive anomaly detection described in the next section.
@@ V-B Bayesian Anomaly Detection
The buffer zone indicates an area around a confined space structure that is neither considered a FOD nor a part of the structure. This buffer zone allows for deviation in the modeling of the structure or objects that are too small to be considered FODs. This buffer zone is specified by a distance threshold, $\epsilon$, offset in the direction of the normal vector of the walls of the reference model, as illustrated as the white area in Fig. 6.
@@
The marginal posterior of a specific inspection node can be obtained from the graph posterior by marginalization. For a Gaussian distribution, it is done by selecting the corresponding entries in $x$ and block from the diagonal of $\Sigma_{x}$ [52]. For instance, the marginal posterior of node $i$ in (68) is
@@
In the $\text{SE}(3)$ formulation, $x_{i}$ is represented by a $4\times 4$ rigid transformation matrix $T_{i}$. The corresponding covariance is a $6\times 6$ matrix with respect to the 6 degrees of freedom of $T_{i}$ on the manifold. The depth point cloud, $\mathcal{P}_{\text{local, i}}$, with a set of points, $p_{\text{local}}\in\mathbb{R}^{3}$, is stored in the local coordinate of its associated pose node, with some independent observation noise $\Sigma_{\text{j,local}}$ for each point $p_{j,\text{local}}$. An example is illustrated in Fig. 7.
Figure 6: Illustration of the half-space Gaussian integrals used for anomaly detection of one point $p_{j}$ in the depth point cloud and the corresponding mapping uncertainty $\Sigma_{j}$. The blue area, $\Omega_{0}$, is the half space wherein the point would be considered to be inside the reference model; and the red area, $\Omega_{1}$ is the half space wherein the point would be considered a foreign object. The black points are from the reference model, $p_{\text{ref}}$ is statistically closest point to $p_{j}$ with normal vector $n_{\text{ref}}$, and $\epsilon$ is the buffer zone threshold. The likelihood of the point $p_{j}$ being sampled from the blue and red areas are obtained by integrating the Gaussian $\mathcal{N}(p_{j},\Sigma_{j})$ over the corresponding halfspaces.
Once the pose graph is optimized, the depth point cloud, $\mathcal{P}_{\text{local,i}}$, can be transformed into the global coordinate with $T_{i}$
@@
with the noise propagation
@@
After transformation, the point cloud can be compared to the reference model $(p_{\text{ref}},n_{\text{ref}})$, where $p_{\text{ref}}$ is the point cloud representation of the reference, and $n_{\text{ref}}$ are the normal vectors of the points, inherited from the plane the points were sampled from. As illustrated in Fig. 6, for each point, $p_{j}$, in the observed point cloud, find the corresponding reference point by
@@
Figure 7: Illustration of the depth point cloud, $\mathcal{P}_{\text{local}}$, associated with an inspection pose node. $T_{i}$ is the $4\times 4$ rigid transformation of the pose node $i$, $\Sigma_{i}$ is the associated covariance matrix, and the small circles, $p_{j}$, are the point cloud samples collected by the depth camera.
Consequently, the likelihood of $p_{j}$ being sampled from the plane formed by $(p_{\text{ref,j}}^{*},n^{*}_{\text{ref,j}})$ or behind with respect to the normal vector, $n_{\text{ref}}$, is calculated by the half space integral
[…]
@@
This can be rewritten by a change of coordinate, resulting in
[…]
$c_{0,j}$ $\triangleq\frac{(n_{\text{ref},j}^{*})^{T}(p_{\text{ref},j}^{*}-p_{j})}{\sqrt{(n_{\text{ref},j}^{*})^{T}\Sigma_{j}n_{\text{ref},j}^{*}}}$
where $c_{0,j}$ is the signed Mahalanobis distance and (75) is an integral of the standard normal distribution, which can be solved by $1-\text{normcdf}(c_{0,j})$. $\mathbb{P}(p_{j}|H_{1})$ is calculated similarly by offsetting $p^{*}_{\text{ref}}$ by $\epsilon$ and integrating from the opposite halfspace.
[…]
$c_{1,j}$ $\triangleq c_{0,j}-\frac{\epsilon}{\sqrt{(n_{\text{ref},j}^{*})^{T}\Sigma_{j}n_{\text{ref},j}^{*}}}.$
@@ V-C Information Measure
Conceptually, the reference point entropy measures the expected surprise of sampling from the reference point $p_{ref,i}$, and the region measure is the maximum amount of surprise the robot can get by sampling from the region. Combining this with the ergodic traversal method, the robot travels to the region with a high probability of being frequently surprised.
@@ VI Hierarchical Ergodic Inspection Algorithm
This section explains how graph ergodic planning and information measure are used within the hierarchical planning framework (as shown in Fig. 1), which we call Hierarchical Ergodic Markov Planning (HEMaP), to inspect a confined structure such as a ballast tank. The complete inspection planning algorithm is described in Algorithm 1. The algorithm requires the initial region of the robot $r[0]$; the region graph $\mathcal{G}$; the replanning horizon $K$; the reference point cloud $\mathcal{P}_{\text{ref}}$; and the initial anomaly entropy $\mu_{p}$. The robot initializes the graph-level time step $k$ to 0. Then, for every $K$ steps, the robot calculates the region-based measure $\mu_{r}$ by summing the anomaly entropy of the structures in each region. Using the new region-based measure $\mu_{r}$, the robot calculates the optimal stochastic transition matrix $P$ according to the REMC optimization problem in (55). Additionally, to prevent an ill-conditioned target distribution with 0-entries, a smoothing factor $\delta$ is added before optimization. The robot samples from this $P$ for the next $K$ steps. After the next region is sampled, the robot picks a waypoint inside the next region according to the structure anomaly entropy in the region (described in the next algorithm), navigates to the waypoint, and updates the anomaly detection via (81). This process is repeated until the inspection is finished, either defined by a maximum number of steps or terminated by a human.
@@
Parameter: number of waypoints per region: $n_{\text{waypoints}}$
Inputs: initial region: $r[0]$, region graph: $\mathcal{G}$, time horizon: $K$, reference point cloud: $\mathcal{P}_{\text{ref}}$, information measure: $\mu_{p}$
Initialize $k\leftarrow 0$
while Inspecting do
     if $k\mod K=0$ then
         Initialize $\mu[n]$
         for $r_{i}\in\mathcal{R}$ do
              for $p_{ref}\in r_{i}$ do
                  $\mu_{r}[i]\leftarrow\mu_{r}[i]+\mu_{p}(p_{ref})$
              end for
         end for
         $\bar{\rho}=\mu_{r}/\sum\mu_{r}$
         $\bar{\rho}=(\bar{\rho}+\delta)/\sum(\bar{\rho}+\delta)$
         $P\leftarrow\text{REMC}(\bar{\rho})$triangleright (55)
     end if
     $r[k+1]\sim\mathbb{P}(R_{k+1}|R_{k}=r[k])$
     for $i$ in $n_{\text{waypoint}}$ do
         $x\leftarrow$ waypoint_placement($r[k+1],\mu_{p}$)
@@
         $\mu_{p}\leftarrow\text{anomaly\_detection}(\mu_{p})$triangleright V-B
     end for
     $k\leftarrow k+1$
end while
Algorithm 2 waypoint_placement
Parameter: number of samples: $n_{\text{sample}}$
Inputs: region: $r$, feature information: $\mu_{p}$
Initialize $x[n_{\text{sample}}],\mu_{x}[n_{\text{sample}}]$
for $i$ in $n_{\text{sample}}$ do
     $x[i]\sim r$
     for $p_{\text{ref}}\in\mathcal{P}_{\text{ref}}$ do
         if $p_{\text{ref}}\in\text{camera\_frustum}(x[i])$ then
              $\mu_{x}[i]\leftarrow\mu_{x}[i]+\mu_{p}(p_{\text{ref}})$
         end if
     end for
end for
$\bar{\mu}_{x}\leftarrow\mu_{x}/\sum\mu_{x}$
$x^{*}\leftarrow\text{random\_choice}(x,p=\bar{\mu}_{x})$
Return $x^{*}$

### v2

@@ IV-E Nonuniform Target Distribution
Selecting the stochastic matrix $P$ to have the smallest, second-largest eigenvalue as in Eq. (17) minimizes the normalized ergodicity deviation in (16) with weighting $Q=\Pi^{-1}$ for a given target distribution $\bar{\rho}$ with no zero entries. Additionally, the optimal stochastic matrix $P$ can be obtained by solving the following convex optimization problem
@@
with the same conditions as in Eq. (29) but with the symmetry $P=P^{T}$ condition replaced by the detailed-balance condition in Eq. (31).
@@
$\max_{\rho_{0}}\frac{\|Q^{1/2}(\mathbb{E}[\hat{\rho}_{w}]-\bar{\rho})\|_{2}}{\|Q^{1/2}\left(\rho_{0}-\bar{\rho}\right)\|_{2}}=\max_{\rho_{0}}\frac{\left\|Q^{1/2}f_{w}\left(P|_{u_{1}^{\perp}}\right)\rho_{0{\perp}}\right\|_{2}}{\|Q^{1/2}\rho_{0\perp}\|_{2}}$ (33)
$\qquad=\max_{\rho_{0}}\frac{\left\|Q^{1/2}f_{w}\left(P|_{u_{1}^{\perp}}\right)Q^{-1/2}Q^{1/2}\rho_{0{\perp}}\right\|_{2}}{\|Q^{1/2}\rho_{0\perp}\|_{2}}$
$\qquad=\max_{\rho_{0}}\frac{\left\|f_{w}\left(Q^{1/2}P|_{u_{1}^{\perp}}Q^{-1/2}\right)Q^{1/2}\rho_{0{\perp}}\right\|_{2}}{\|Q^{1/2}\rho_{0\perp}\|_{2}}$
$\qquad=\max_{\rho_{0}}\frac{\left\|f_{w}\left(\tilde{P}|_{u_{1}^{\perp}}\right)\tilde{\rho}_{0{\perp}}\right\|_{2}}{\|\tilde{\rho}_{0\perp}\|_{2}}$
@@
where $\Pi=\text{diag}(\bar{\rho})$ is invertible and its square root exists since $\bar{\rho}$ is strictly positive. Note that $\tilde{P}$ is symmetric since the detailed balance condition in (31), results in
$\Pi^{-1/2}P\Pi^{1/2}$ $=\Pi^{1/2}P^{T}\Pi^{-1/2}$ (35)
i.e., $\tilde{P}=\tilde{P}^{T}$, which also implies that the restriction of $\tilde{P}$ to the subspace orthogonal to $u_{1}$, i.e., $\tilde{P}|_{u_{1}^{\perp}}$ is symmetric, since all the eigen-components, $\lambda_{i}v_{i}u_{i}^{T}$, of $\tilde{P}$ are symmetric. Then, following arguments as in Eqs. (23)- (27) in the proof ofTherorem 2,
@@ IV-F Convex Bounding for General Markov Chain
The detailed balance condition in the previous subsection can be restrictive for ergodicity. An obvious example is when transitions are allowed in only one direction between some regions, because the detailed balance condition leads to zero probability even for the available transition. Thus, in such one-way transition cases, the detailed balance condition results in a larger normalized ergodicity deviation compared to the case without the condition, as illustrated in Fig. 4.
@@
The detailed balance requirement can be relaxed for general directed graphs (e.g., with potential one-way transitions), by optimizing an upper bound on the normalized ergodicity deviation, as shown in the theorem below.
@@
The following corollary states the convex optimization problem for Theorem 3, which is referred in the paper as REMC.
Corollary 3.1 (Rapidly Ergodic Markov Chain (REMC)).
The optimal $P$ for the upper-bound in Theorem 3 can be obtained by solving the convex optimization problem
@@ IV-G Connection to Fastest Mixing Markov Chain
This is equivalent to the fastest mixing Markov chain (FMMC) formulation, where the stochastic matrix $P$ is selected such that the SLEM, $|\lambda_{2}|$, is minimized. The FMMC formulation upper-bounds the REMC by ensuring fast convergence on the discounted time average since,
@@
The difference in ergodicity achieved with the FMMC formulation that minimizes the SLEM and the proposed REMC approach that minimizes the second largest eigenvalue is illustrated in the following 2-node example. An additional example with a 3-node graph is illustrated in Appendix B.
@@ V Estimation Framework
We now discuss the estimation module in Fig. 1 that comprises SLAM and anomaly detection. In addition to the occupancy grid map and pose estimates for navigation, the estimator also provides the measure $\mu$ (defined in (3)) to the REMC method. Specifically, this measure is selected as information entropy using a Bayesian anomaly (FOD) detection method. The core idea of detecting FODs is to evaluate the hypothesis that the difference between the local (spatial) point cloud given by the robot’s depth camera and the reference model of the corresponding section of the confined space is significant. Consequently, the FOD detection method uses the marginal posterior of our SLAM formulation described in the following Section.
In short, the SLAM module (Section V-A) provides optimal estimates of the robot’s poses and map features along with their uncertainties. These estimates are then used to transform the local spatial point cloud, sampled by the RGB-D camera, into the global coordinate system, which subsequently help with anomaly detection (Section.V-B). Finally, the information metric $\mu_{p}$ is generated from the confidence level of the anomaly detection. A flow chart outlining the steps is shown in Fig. 6.
Figure 6: Flow chart of the estimation module. Hierarchical SLAM processes the sensor information to generate a Gaussian posterior of the pose, $(\mathcal{T},\Sigma_{\mathcal{T}})$, detailed implementation of which is discussed in Section V-A. The anomaly detector compares the local point cloud from the RGB-D camera $\mathcal{P}_{\text{local}}$ with the reference point cloud $\mathcal{P}_{\text{ref}}$ using the posterior to generate the information map $\mu_{p}$.
@@ V-A Hierarchical SLAM
The SLAM module serves two purposes in our inspection framework. First, it provides a probabilistic mapping of the environment in the form of a 3D point cloud of the structures that is used for FOD detection via Bayesian inference. Second, it estimates the robot pose (location and orientation with respect to the map origin), enabling navigation within the space.To accommodate non-linear robot motion and sensor models, we use the Graph-SLAM algorithm. In this method, a factor graph is constructed during run-time based on the collected observations and odometry, with nodes representing the robot’s poses and map features.
Specifically, the poses are calculated on the $\text{SE}(3)$ manifold, i.e., a set of 3D rigid transformation matrices, according to the formulation in [51], and are optimized by the Gauss–Newton algorithm described in [52]. We modify it as a hierarchical SLAM, similar to [53] and [54]. Instead of maintaining the full resolution factor graph and performing hierarchical optimization, marginalization is done as an intermediate EKF-SLAM, and only the sparse factor graph is maintained with the posterior from the EKF-SLAM as the factor between the nodes. In particular, EKF-SLAM is reset and a new pose node is created based on the distance traveled from the most recent pose node. Special nodes, called inspection nodes, are created once the robot reaches the waypoint specified by the waypoint placement method. The corresponding depth point clouds, i.e., the point clouds generated by the depth images taken by the robot at the waypoints, are also stored at these inspection nodes. Furthermore, nodes that fall outside the current estimation horizon are pruned and replaced by a factor node through marginalization. Once the Graph-SLAM is optimized, the estimated poses of all the unpruned nodes, $\mathcal{T}\triangleq\{T_{1},T_{2},\cdots T_{i}\cdots\}$, and the corresponding covariance matrices, $\Sigma_{\mathcal{T}}\triangleq\text{Var}(\mathcal{T})$, are used for recursive anomaly detection described in the next Section.
@@ V-B Bayesian Anomaly Detection
The buffer zone indicates an area around a confined space structure that is neither considered a FOD nor a part of the structure. This buffer zone allows for deviation in the modeling of the structure or objects that are too small to be considered FODs. This buffer zone is specified by a distance threshold, $d_{\text{buffer}}$, offset in the direction of the normal vector of the walls of the reference model, as illustrated in Fig. 7.
@@
The marginal posterior of a particular inspection node can be obtained from the graph posterior by marginalization. For a Gaussian distribution, it is done by selecting the corresponding entries in $x$ and blocks from the diagonal of $\Sigma_{x}$ [55]. For instance, the marginal posterior of node $i$ in (60) is
@@
In the $\text{SE}(3)$ formulation, $x_{i}$ is represented by a $4\times 4$ rigid transformation matrix $T_{i}$. The corresponding covariance is a $6\times 6$ matrix with respect to the 6 degrees of freedom of $T_{i}$ on the manifold. The local point cloud, $\mathcal{P}_{\text{local},i}$, with a set of points, $p_{\text{local}}\in\mathbb{R}^{3}$, is stored in the local coordinate of its associated pose node, with some independent observation noise $\Sigma_{j,\text{,local}}$ for each point $p_{j,\text{local}}$. An example is illustrated in Fig. 7(a).
Once the pose graph is optimized, the local point cloud, $\mathcal{P}_{\text{local},i}$, is transformed into the global coordinate with $T_{i}$
@@
with the noise propagated as
@@
After transformation, the point cloud is compared with the reference model $(p_{\text{ref}},n_{\text{ref}})$, where $p_{\text{ref}}$ is the representation of the point cloud and $n_{\text{ref}}$ are the normal vectors of the points that are inherited from the plane the points are sampled from. As illustrated in Fig. 7(b), for each point, $p_{j}$, in the observed point cloud, we find the corresponding reference point as
@@
Figure 7: SLAM-based anomaly detection. (a) shows an illustration of a local point cloud $\mathcal{P}_{\text{local}}$ of an inspection pose node. $T_{i}$ is the $4\times 4$ rigid transformation of the pose node $i$, $\Sigma_{i}$ is the associated covariance matrix, and the small circles, $p_{j}$, are the point cloud samples collected by the depth camera. (b) illustrates the associated FOD detection on a single point $p_{j}\in\mathcal{P}_{\text{local}}$ with the uncertainty $\Sigma_{j}$ propagated from (a). The blue area, $\Omega_{0}$, is the half space wherein the point would be considered to be inside the reference mesh; and the red area, $\Omega_{1}$ is the half space wherein the point would be considered a foreign object. The black points are from the reference model, $p_{\text{ref}}$ is statistically closest point to $p_{j}$ with normal vector $n_{\text{ref}}$, and $d_{\text{buffer}}$ is the buffer zone threshold. The likelihood of the point $p_{j}$ being sampled from the blue and red areas are obtained by integrating the Gaussian $\mathcal{N}(p_{j},\Sigma_{j})$ over the corresponding halfspaces.
Consequently, the likelihood of $p_{j}$ being sampled from the plane formed by $(p_{\text{ref},j}^{*},n^{*}_{\text{ref},j})$ or behind with respect to the normal vector, $n_{\text{ref}}$, is calculated by the half-space integral
$\mathbb{P}(p_{j}|H_{0})$ $=\int_{\Omega_{0}}\mathcal{N}(p_{j},\Sigma_{j})dx$ (65)
@@
This is calculated by the normal cumulative distribution function (cdf) as
$\mathbb{P}(p_{j}|H_{0})$ $=1-\text{normcdf}(c_{0,j})$ (66)
where $c_{0,j}$ is the signed Mahalanobis distance
$c_{0,j}$ $\triangleq\frac{(n_{\text{ref},j}^{*})^{T}(p_{\text{ref},j}^{*}-p_{j})}{\sqrt{(n_{\text{ref},j}^{*})^{T}\Sigma_{j}n_{\text{ref},j}^{*}}}.$ (67)
$\mathbb{P}(p_{j}|H_{1})$ is calculated similarly by offsetting $p^{*}_{\text{ref}}$ by $d_{\text{buffer}}$ and integrating from the opposite halfspace by $\mathbb{P}(p_{j}|H_{1})=\text{normcdf}(c_{1,j})$, with
$c_{1,j}$ $\triangleq c_{0,j}-\frac{d_{\text{buffer}}}{\sqrt{(n_{\text{ref},j}^{*})^{T}\Sigma_{j}n_{\text{ref},j}^{*}}}.$ (68)
@@ V-C Information Measure
Conceptually, the reference point entropy measures the expected surprise of sampling from the reference point $p_{ref,i}$, and the region measure is the maximum amount of surprise the robot can get by sampling from the region. Combining this with the REMC, the robot travels to the region with a high probability of being frequently surprised.
@@ VI Hierarchical Ergodic Inspection Algorithm
This section explains how graph ergodic planning and information measure are used within the hierarchical planning framework (as shown in Fig. 1), which we call hierarchical ergodic Markov planner (HEMaP), to inspect a confined structure such as a ballast tank. The complete inspection planning algorithm is described in Algorithm 1. The algorithm requires the initial region of the robot $r[0]$; the region graph $\mathcal{G}$; the replanning horizon $K$; the reference point cloud $\mathcal{P}_{\text{ref}}$; and the initial anomaly entropy $\mu_{p}$. The robot initializes the graph-level time step $k$ to 0. Then, for every $K$ steps, the robot calculates the region-based measure $\mu_{r}$ by summing the anomaly entropy of the structures in each region. Using the new region-based measure $\mu_{r}$, the robot calculates the optimal stochastic transition matrix $P$ according to the REMC optimization problem in (47). Additionally, to prevent an ill-conditioned target distribution with 0-entries, a smoothing factor $\delta$ is added before optimization. The robot samples from this $P$ for the next $K$ steps. After the next region is sampled, the robot picks a waypoint inside the next region according to the structure anomaly entropy in the region (described in the next algorithm), navigates to the waypoint, and updates the anomaly detection via (74). This process is repeated until the inspection is finished, either defined by a maximum number of steps or terminated by a human.
@@
1:Parameter: number of waypoints per region: $n_{\text{waypoints}}$
2:Inputs: initial region: $r[0]$, region graph: $\mathcal{G}$, time horizon: $K$, reference point cloud: $\mathcal{P}_{\text{ref}}$, information measure: $\mu_{p}$
3:Initialize $k\leftarrow 0$
4:while Inspecting do
5:  if $k\mod K=0$ then
6:   Initialize $\mu[n]$
7:   for $r_{i}\in\mathcal{R}$ do
8:     for $p_{ref}\in r_{i}$ do
9:      $\mu_{r}[i]\leftarrow\mu_{r}[i]+\mu_{p}(p_{ref})$
10:     end for
11:   end for
12:   $\bar{\rho}=\mu_{r}/\sum\mu_{r}$
13:   $\bar{\rho}=(\bar{\rho}+\delta)/\sum(\bar{\rho}+\delta)$
14:   $P\leftarrow\text{REMC}(\bar{\rho})$ $\triangleright$ (47)
15:  end if
16:  $r[k+1]\sim\mathbb{P}(R_{k+1}|R_{k}=r[k])$
17:  for $i$ in $n_{\text{waypoint}}$ do
18:   $x\leftarrow$ waypoint_placement($r[k+1],\mu_{p}$)
19:   $\text{navigate\_to}(x)$
20:   $\mu_{p}\leftarrow\text{anomaly\_detection}(\mu_{p})$ $\triangleright$ V-B
21:  end for
22:  $k\leftarrow k+1$
23:end while
Algorithm 2 waypoint_placement
1:Parameter: number of samples: $n_{\text{sample}}$
2:Inputs: region: $r$, feature information: $\mu_{p}$
3:Initialize $x[n_{\text{sample}}],\mu_{x}[n_{\text{sample}}]$
4:for $i$ in $n_{\text{sample}}$ do
5:  $x[i]\sim\text{uniform}(r)$ $\triangleright$ uniformly sample a pose from the region
6:  for $p_{\text{ref}}\in\mathcal{P}_{\text{ref}}$ do
7:   if $p_{\text{ref}}\in\text{camera\_frustum}(x[i])$ then
8:     $\mu_{x}[i]\leftarrow\mu_{x}[i]+\mu_{p}(p_{\text{ref}})$
9:   end if
10:  end for
11:end for
@@
[…]
14:Return $x^{*}$